A quality-conscious disk manufacturer wishes to know the fraction of disks his company makes which are defective. Step 1 of 2 : Suppose a sample of 707 floppy disks is drawn. Of these disks, 77 were defective. Using the data, estimate the proportion of disks which are defective. Enter your answer as a fraction or a decimal number rounded to three decimal places.
Answer #1

Solution :

Given that,

n = 707

x =77

Point estimate = sample proportion = \hat p = x / n =77 / 707=0.109

answer =0.109